# TokenCLIP: Token-wise Prompt Learning for Zero-shot Anomaly Detection

### v3

@@ 1 Introduction
Foundation Models (FMs) (radford2021learning; kirillov2023segment; qwen2025qwen25technicalreport) have shown the potential to generalize to unseen class semantics and domains. This breakthrough has driven the rapid development of downstream tasks that explore zero-shot capabilities by adapting FMs (pang2021deep; zhou2022learning; khattak2023maple; jeong2023winclip; zhou2023anomalyclip; gu2024anomalygpt). Anomaly detection has also followed this trend, evolving from specialized models toward more generalized detection frameworks (jeong2023winclip; zhou2023anomalyclip; chen2023zero; Cao2024AdaCLIPAC; qu2025bayesianpromptflowlearning; jiang2025mmadcomprehensivebenchmarkmultimodal; zhou2024pointad; gao2025adaptclipadaptingclipuniversal).
A prominent line of research in this area involves adapting CLIP for zero-shot anomaly detection. These methods typically project either learnable (zhou2023anomalyclip; gu2024anomalygpt) or handcrafted text prompts (jeong2023winclip; chen2023zero) into a shared embedding space and align them with diverse visual features, capturing both global and local abnormalities. Despite their appeal, these approaches rely on a single textual space to indiscriminately align with all visual tokens, whether detecting a crack on a carpet or a tumor in a brain scan. As shown in Figure LABEL:fig:1_1, this coarse alignment makes it difficult for the detection model to capture generalized anomaly semantics, as the token-agnostic textual space is forced to make a tradeoff between diverse semantic tokens. As a result, the model tends to favor common anomalies while compromising the rare anomaly semantics.
@@ 2 Related Work
ZSAD is an emerging field that aims to detect anomalies in unseen object categories and even across domains (esmaeilpour2022zero; li2023zero; gao2025metauasuniversalanomalysegmentation). Some methods, such as VAND (chen2023zero) and AdaCLIP (Cao2024AdaCLIPAC), focus on adapting CLIP’s visual encoder to capture anomaly semantics. However, these approaches rely heavily on human-crafted text prompts to represent normality and abnormality (jeong2023winclip). Another branch, including methods like AnomalyCLIP (zhou2023anomalyclip), FiLo (Gu2024FiLoZA) takes the opposite approach: rather than adapting the visual space, they learn text prompts to adapt the textual space for modeling anomaly semantics. Additionally, methods such as AACLIP (Ma2025AACLIPEZ), BayesCLIP (qu2025bayesianpromptflowlearning), and AdaptCLIP (gao2025adaptclipadaptingclipuniversal) aim to adapt both the visual and textual spaces for improved performance. Despite these advances, most existing methods rely on a single textual space to simultaneously align with diverse visual patch tokens. FAPrompt (zhu2024finegrainedabnormalitypromptlearning) ensembles multiple prompts to capture more abnormal patterns, but remains at the image level. In contrast, TokenCLIP introduces a dynamic alignment that provides token-level supervision by assigning each visual patch token to a semantics-aware weighted combination of textual subspaces.
@@
Prompt learning was proposed to efficiently adapt CLIP for more accurate image classification with minimal computational overhead (zhou2022learning). AnomalyCLIP (zhou2023anomalyclip) extends prompt learning to zero-shot anomaly detection, aiming to capture both local and global semantics for anomaly classification and localization Gu2024FiLoZA. They typically model a token-agnostic textual space and overlook the semantic differences among local regions. In contrast, TokenCLIP introduces an orthogonal textual subspace and adaptively combines these subspaces to align each token according to its visual semantics. We formulate this dynamic alignment between visual regions and textual prompts as an OT problem. While PLOT (chen2023plotpromptlearningoptimal) also incorporates OT for prompt learning, it is primarily designed for image-level classification. In contrast, we leverage OT for pixel-level anomaly segmentation and image-level anomaly detection, achieving fine-grained and spatially aware alignment.
@@
OT has emerged as a powerful framework for comparing probability distributions, with wide applications in computer vision (Villani2008OptimalTO). The entropic regularized OT proposed by Cuturi (Cuturi2013SinkhornDL) significantly improved computational efficiency via the Sinkhorn algorithm, making OT feasible for large-scale learning. Subsequent works (genevay2016stochastic; 8641476) have extended OT to stochastic, unbalanced, and mini-batch scenarios. In multi-modal learning, OT has been employed for fine-grained alignment between modalities (e.g., image patches and text tokens), enabling interpretable and structure-aware correspondence. In contrast to prior work chen2023plotpromptlearningoptimal, we are the first to introduce OT for fine-grained anomaly semantics learning, particularly in local visual anomaly detection. We observe that standard OT often results in overly dense transport plans and propose a topK selection mechanism to enforce cleaner, more discriminative alignments between visual and textual spaces.
@@ 4 Experiments
We evaluate TokenCLIP in ZSAD through large-scale experiments. The evaluation covers two distinct domains: industrial inspection and medical diagnosis. In the industrial domain, we evaluate seven benchmarks: MVTec AD (bergmann2019mvtec), VisA (zou2022spot), MPDD (jezek2021deep), BTAD (mishra2021vt), SDD (tabernik2020segmentation), DAGM (wieler2007weakly), and DTD-Synthetic (aota2023zero). These datasets span various manufactured objects and defect types. In the medical domain, we assess tasks including skin lesion detection (ISIC (gutman2016skin)), colon polyp segmentation (CVC-ClinicDB (bernal2015wm), CVC-ColonDB (tajbakhsh2015automated), Kvasir (jha2020kvasir), Endo (hicks2021endotect)), and brain anomaly detection (HeadCT, BrainMRI (salehi2021multiresolution), Br35H (br35h)). Detailed dataset, implementation, and evaluation metric are provided in Appendices AB, and C.
@@
Image-level
(AUROC, AP)
MVTec AD (88.8, 94.8) (91.8, 96.5) (86.1, 93.5) (89.6, -) (91.5, 96.2) (90.8, 94.9) (93.5, 96.7)
@@
Pixel-level
(AUROC, PRO)
MVTec AD (33.3, 6.7) (85.1, 64.6) (87.6, 44.0) (90.3, -) (91.1, 81.4) (90.6, 81.6) (92.2, 87.9)
@@ 4.1 Main results
Image-level
(AUROC, AP)
HeadCT (78.4, 78.8) (81.8, 80.2) (89.1, 89.4) (91.5, -) (93.4, 91.6) (93.9, 92.6) (96.0, 95.3)
@@
Pixel-level
(AUROC, PRO)
ISIC (51.7, 15.9) (83.3, 55.1) (89.4, 77.2) (88.3, -) (89.7, 78.4) (90.7, 80.3) (91.6, 83.4)
@@ 4.2 Result analysis
Here, we discuss the contributions of OT and orthogonality regularization to textual specialization. To this end, we visualize the subspace assignment frequency on background and foreground regions across different categories. The first row shows the results of TokenCLIP. The green subspace mainly focuses on object foregrounds (FG), while the blue and red subspaces focus on backgrounds (BG). This demonstrates that TokenCLIP successfully achieves subspace specialization, consistent with the observation in Figure LABEL:fig:2_2. The middle row shows TokenCLIP without OT. In this case, no clear textual specialization is observed. For example, in the Toothbrush category, the green subspace dominates both FG and BG assignments, while in Hazelnut it is rarely used. This indicates that OT plays a dominant role in enabling fine-grained semantic specialization. The bottom row shows TokenCLIP without orthogonality regularization. The red subspace dominates the FG regions, while the blue and green subspaces are mainly assigned to the BG regions. The specialization still emerges even without orthogonality regularization, but the extent of specialization is weaker than that of the full TokenCLIP model. We additionally visualize the token-level assignments of TokenCLIP without orthogonality regularization in Figure 14.
@@
Methods
Training
per epoch
Inference
Time
FPS
Peak GPU
Memory
MVTec AD VisA
@@ 5 Ablation study
Subspace
number $Q$
MVTec AD VisA
@@
topK
MVTec AD VisA
@@ Appendix B Baselines
Our approach replaces conventional static alignment with a dynamic alignment strategy and can serve as a plug-in for CLIP-based anomaly detection methods. Given the rapid evolution of the field, we compare TokenCLIP against several representative methods that adopt static alignment: CoOp (zhou2022learning), WinCLIP (jeong2023winclip), VAND (chen2023zero), AnomalyCLIP (zhou2023anomalyclip), AdaCLIP (Cao2024AdaCLIPAC), and FAprompt (zhu2024finegrainedabnormalitypromptlearning).
  • CoOp (IJCV 2022) (zhou2022learning): A prompt optimization method that replaces fixed text templates with learnable embeddings. Following zhou2023anomalyclip, we construct prompts by inserting tokens representing normal or anomalous conditions before the class name. Specifically, the templates take the form $V_{1}][V_{2}]...[V_{N}]\,\texttt{normal}\,\texttt{[cls]}$ and $V_{1}][V_{2}]...[V_{N}]\,\texttt{anomalous}\,\texttt{[cls]}$.
  • WinCLIP (CVPR 2023) (jeong2023winclip): They leverage a comprehensive set of handcrafted prompts tailored to anomaly scenarios. It introduces a window-based scaling mechanism to enhance anomaly localization. We reproduce the experimental setup as reported in the original publication for consistency.
  • VAND (ARXIV 2023) (chen2023zero): This method advances prompt design by incorporating learnable linear projections, allowing it to better capture fine-grained visual semantics. We adopt the original implementation and settings to align with the authors’ reported results.
  • AdaCLIP (ECCV 2024) (Cao2024AdaCLIPAC): AdaCLIP retains handcrafted textual prompts but focuses on adapting the visual embedding space to improve anomaly detection.
  • AnomalyCLIP (ICLR 2024) (zhou2022learning): AnomalyCLIP introduces object-agnostic prompt learning and achieves promising generalization across different types of anomalies. It also adapts the textual and visual spaces for better detection performance.
  • FAPromt (ICCV 2025) (zhu2024finegrainedabnormalitypromptlearning): FAPromt proposes to use multiple learnable text prompts to learn complementary and decomposed abnormality. prompts
@@ Appendix C Evaluation Setting and Metrics
We evaluate image-level detection using AUROC and Average Precision (AP). For pixel-level segmentation, we report AUROC and AUPRO. While AUROC reflects overall pixel-wise discrimination, AUPRO emphasizes the quality of region-level anomaly localization. All results are averaged over five independent runs for robustness. Following the evaluation setting (zhou2023anomalyclip), MVTec AD serves as the auxiliary training set when testing on other datasets. Conversely, VisA is used for training when evaluating on MVTec AD. Final results are computed by averaging over all sub-datasets within each benchmark.
@@ Appendix D Hyperparameter Ablation
$\epsilon$
MVTec AD VisA

### v4

@@ 1 Introduction
Foundation Models (FMs) (Radford et al., 2021; Kirillov et al., 2023; Qwen et al., 2025) have shown the potential to generalize to unseen class semantics and domains. This breakthrough has driven the rapid development of downstream tasks that explore zero-shot capabilities by adapting FMs (Pang et al., 2021; Zhou et al., 2022; Khattak et al., 2023; Jeong et al., 2023; Zhou et al., 2024a; Gu et al., 2024b). Anomaly detection has also followed this trend, evolving from specialized models toward more generalized detection frameworks (Jeong et al., 2023; Zhou et al., 2024a; Chen et al., 2023b; Cao et al., 2024; Qu et al., 2025; Jiang et al., 2025; Zhou et al., 2024b; Gao et al., 2025).
A prominent line of research in this area involves adapting CLIP for zero-shot anomaly detection. These methods typically project either learnable (Zhou et al., 2024a; Gu et al., 2024b) or handcrafted text prompts (Jeong et al., 2023; Chen et al., 2023b) into a shared embedding space and align them with diverse visual features, capturing both global and local abnormalities. Despite their appeal, these approaches rely on a single textual space to indiscriminately align with all visual tokens, whether detecting a crack on a carpet or a tumor in a brain scan. As shown in Figure LABEL:fig:1_1, this coarse alignment makes it difficult for the detection model to capture generalized anomaly semantics, as the token-agnostic textual space is forced to make a tradeoff between diverse semantic tokens. As a result, the model tends to favor common anomalies while compromising the rare anomaly semantics.
@@ 2 Related Work
ZSAD is an emerging field that aims to detect anomalies in unseen object categories and even across domains (Esmaeilpour et al., 2022; Li et al., 2023; Gao, 2024). Some methods, such as VAND (Chen et al., 2023b) and AdaCLIP (Cao et al., 2024), focus on adapting CLIP’s visual encoder to capture anomaly semantics. However, these approaches rely heavily on human-crafted text prompts to represent normality and abnormality (Jeong et al., 2023). Another branch, including methods like AnomalyCLIP (Zhou et al., 2024a), FiLo (Gu et al., 2024a) takes the opposite approach: rather than adapting the visual space, they learn text prompts to adapt the textual space for modeling anomaly semantics. Additionally, methods such as AACLIP (Ma et al., 2025), BayesCLIP (Qu et al., 2025), and AdaptCLIP (Gao et al., 2025) aim to adapt both the visual and textual spaces for improved performance. Despite these advances, most existing methods rely on a single textual space to simultaneously align with diverse visual patch tokens. FAPrompt (ZHU et al., 2025) ensembles multiple prompts to capture more abnormal patterns, but remains at the image level. In contrast, TokenCLIP introduces a dynamic alignment that provides token-level supervision by assigning each visual patch token to a semantics-aware weighted combination of textual subspaces.
@@
Prompt learning was proposed to efficiently adapt CLIP for more accurate image classification with minimal computational overhead (Zhou et al., 2022). AnomalyCLIP (Zhou et al., 2024a) extends prompt learning to zero-shot anomaly detection, aiming to capture both local and global semantics for anomaly classification and localization Gu et al. (2024a). They typically model a token-agnostic textual space and overlook the semantic differences among local regions. In contrast, TokenCLIP introduces an orthogonal textual subspace and adaptively combines these subspaces to align each token according to its visual semantics. We formulate this dynamic alignment between visual regions and textual prompts as an OT problem. While PLOT (Chen et al., 2023a) also incorporates OT for prompt learning, it is primarily designed for image-level classification. In contrast, we leverage OT for pixel-level anomaly segmentation and image-level anomaly detection, achieving fine-grained and spatially aware alignment.
@@
OT has emerged as a powerful framework for comparing probability distributions, with wide applications in computer vision (Villani and others, 2008). The entropic regularized OT proposed by Cuturi (Cuturi, 2013) significantly improved computational efficiency via the Sinkhorn algorithm, making OT feasible for large-scale learning. Subsequent works (Genevay et al., 2016; Peyré and Cuturi, 2019) have extended OT to stochastic, unbalanced, and mini-batch scenarios. In multi-modal learning, OT has been employed for fine-grained alignment between modalities (e.g., image patches and text tokens), enabling interpretable and structure-aware correspondence. In contrast to prior work Chen et al. (2023a), we are the first to introduce OT for fine-grained anomaly semantics learning, particularly in local visual anomaly detection. We observe that standard OT often results in overly dense transport plans and propose a topK selection mechanism to enforce cleaner, more discriminative alignments between visual and textual spaces.
@@ 4 Experiments
We evaluate TokenCLIP in ZSAD through large-scale experiments. The evaluation covers two distinct domains: industrial inspection and medical diagnosis. In the industrial domain, we evaluate seven benchmarks: MVTec AD (Bergmann et al., 2019), VisA (Zou et al., 2022), MPDD (Jezek et al., 2021), BTAD (Mishra et al., 2021), SDD (Tabernik et al., 2020), DAGM (Wieler and Hahn, 2007), and DTD-Synthetic (Aota et al., 2023). These datasets span various manufactured objects and defect types. In the medical domain, we assess tasks including skin lesion detection (ISIC (Gutman et al., 2016)), colon polyp segmentation (CVC-ClinicDB (Bernal et al., 2015), CVC-ColonDB (Tajbakhsh et al., 2015), Kvasir (Jha et al., 2020), Endo (Hicks et al., 2021)), and brain anomaly detection (HeadCT, BrainMRI (Salehi et al., 2021), Br35H (Hamada., 2020)). Detailed dataset, implementation, and evaluation metric are provided in Appendices AB, and C.
@@
Image-level (AUROC, AP) MVTec AD (88.8, 94.8) (91.8, 96.5) (86.1, 93.5) (89.6, -) (91.5, 96.2) (90.8, 94.9) (93.5, 96.7)
@@
Pixel-level (AUROC, PRO) MVTec AD (33.3, 6.7) (85.1, 64.6) (87.6, 44.0) (90.3, -) (91.1, 81.4) (90.6, 81.6) (92.2, 87.9)
@@ 4.1 Main results
Image-level (AUROC, AP) HeadCT (78.4, 78.8) (81.8, 80.2) (89.1, 89.4) (91.5, -) (93.4, 91.6) (93.9, 92.6) (96.0, 95.3)
@@
Pixel-level (AUROC, PRO) ISIC (51.7, 15.9) (83.3, 55.1) (89.4, 77.2) (88.3, -) (89.7, 78.4) (90.7, 80.3) (91.6, 83.4)
@@ 4.2 Result analysis
Here, we discuss the contributions of OT and orthogonality regularization to textual specialization. To this end, Figure 14(a) presents the subspace assignment frequency on background and foreground regions across different categories. The first row shows the results of TokenCLIP. The green subspace mainly focuses on object foregrounds (FG), while the blue and red subspaces focus on backgrounds (BG). This demonstrates that TokenCLIP successfully achieves subspace specialization, consistent with the observation in Figure LABEL:fig:2_2. The middle row shows TokenCLIP without OT. In this case, no clear textual specialization is observed. For example, in the Toothbrush category, the green subspace dominates both FG and BG assignments, while in Hazelnut it is rarely used. This indicates that OT plays a dominant role in enabling fine-grained semantic specialization. The bottom row shows TokenCLIP without orthogonality regularization. The red subspace dominates the FG regions, while the blue and green subspaces are mainly assigned to the BG regions. The specialization still emerges even without orthogonality regularization, but the extent of specialization is weaker than that of the full TokenCLIP model. We additionally visualize the token-level assignments of TokenCLIP without orthogonality regularization in Figure 14.
@@
Methods Training per epoch Inference Time FPS Peak GPU Memory MVTec AD VisA
@@ 5 Ablation study
Subspace number $Q$ MVTec AD VisA
@@
topK MVTec AD VisA
@@ Appendix B Baselines
Our approach replaces conventional static alignment with a dynamic alignment strategy and can serve as a plug-in for CLIP-based anomaly detection methods. Given the rapid evolution of the field, we compare TokenCLIP against several representative methods that adopt static alignment: CoOp (Zhou et al., 2022), WinCLIP (Jeong et al., 2023), VAND (Chen et al., 2023b), AnomalyCLIP (Zhou et al., 2024a), AdaCLIP (Cao et al., 2024), and FAprompt (ZHU et al., 2025).
  • CoOp (IJCV 2022) (Zhou et al., 2022): A prompt optimization method that replaces fixed text templates with learnable embeddings. Following Zhou et al. (2024a), we construct prompts by inserting tokens representing normal or anomalous conditions before the class name. Specifically, the templates take the form $V_{1}][V_{2}]...[V_{N}]\,\texttt{normal}\,\texttt{[cls]}$ and $V_{1}][V_{2}]...[V_{N}]\,\texttt{anomalous}\,\texttt{[cls]}$.
  • WinCLIP (CVPR 2023) (Jeong et al., 2023): They leverage a comprehensive set of handcrafted prompts tailored to anomaly scenarios. It introduces a window-based scaling mechanism to enhance anomaly localization. We reproduce the experimental setup as reported in the original publication for consistency.
  • VAND (ARXIV 2023) (Chen et al., 2023b): This method advances prompt design by incorporating learnable linear projections, allowing it to better capture fine-grained visual semantics. We adopt the original implementation and settings to align with the authors’ reported results.
  • AdaCLIP (ECCV 2024) (Cao et al., 2024): AdaCLIP retains handcrafted textual prompts but focuses on adapting the visual embedding space to improve anomaly detection.
  • AnomalyCLIP (ICLR 2024) (Zhou et al., 2022): AnomalyCLIP introduces object-agnostic prompt learning and achieves promising generalization across different types of anomalies. It also adapts the textual and visual spaces for better detection performance.
  • FAPromt (ICCV 2025) (ZHU et al., 2025): FAPromt proposes to use multiple learnable text prompts to learn complementary and decomposed abnormality. prompts
@@ Appendix C Evaluation Setting and Metrics
We evaluate image-level detection using AUROC and Average Precision (AP). For pixel-level segmentation, we report AUROC and AUPRO. While AUROC reflects overall pixel-wise discrimination, AUPRO emphasizes the quality of region-level anomaly localization. All results are averaged over five independent runs for robustness. Following the evaluation setting (Zhou et al., 2024a), MVTec AD serves as the auxiliary training set when testing on other datasets. Conversely, VisA is used for training when evaluating on MVTec AD. Final results are computed by averaging over all sub-datasets within each benchmark.
@@ Appendix D Hyperparameter Ablation
$\epsilon$ MVTec AD VisA